A chemist wants to make 250 ml of 30% acid solution by mixing a 20% acid solution with a 60% acid solution. determine how much acid solution should be used as follows. x represents the amount of 20% acid solution. y represents the amount of 60% acid solution.
write an equation relating x and y to the desired amount of acid.

Answer by checkley77(12844)

x+y=250 or y=250-x
.20x+.60y=.30*250
.20x+.60(250-x)=75
.20+150-.60x=75
-.40x=75=150
-.40x=75
x=75/.40
x=187.5 ml. of 20% mix is used.
250-187.5=62.5 ml. of 60% mix is uded.
Proof:
.20*187.5+.60*62.5=75
37.5+37.5=75
75=75